# Probing forced responses and causal mechanisms in large-scale climate dynamics with reduced-order neural models

### v3

@@ IV.2.3 Interpretability and understanding: ensemble mean and variance responses
In the previous section, we showed that the reduced-order neural emulator can reconstruct the net radiative flux at the top of the atmosphere (TOA) from prescribed changes in the sea surface temperature (SST) field. This amounts to the mapping $\mathbf{x}_{\text{SST}}^{\text{f}}(t)\rightarrow x_{\text{TOA}}(t+1)$, where $\mathbf{x}_{\text{SST}}(t)\in\mathbb{R}^{n}$, $x_{\text{TOA}}(t)\in\mathbb{R}$, and $x_{\text{TOA}}(0)=0$. We now ask the following question: What are the SST patterns that contribute the most to this skillfull prediction? Addressing this question requires moving beyond prediction to identify the causal influence of individual SST components on radiative flux. We frame the problem in terms of linear response theory, focusing on the direct link $x_{j,\text{SST}}(t)\rightarrow x_{\text{TOA}}(t+1)$, where $x_{j,\text{SST}}(t)$ denotes the $j$-th SST component.
Response in ensemble mean.
To quantify this link, we perturb the emulator in Eq. (11) following the procedure in Appendix A. Briefly, at $t=0$, we apply a small impulse $\Delta_{j}=10^{-2}\sigma_{j}$ to the $j$-th SST component, with $\sigma_{j}$ its stationary standard deviation. We then compute the ensemble-mean response of the net TOA flux at $t=1$, denoted $\delta\langle x_{\text{TOA}}(1)\rangle$. This quantity, referred to as $R^{\mu}_{\text{TOA},j}(1)$, represents the ensemble-mean response ($\mu$ indicates the mean) and is estimated over $10^{6}$ initial conditions to ensure robust statistics
Response in ensemble variance.
While the ensemble-mean response is informative, it captures only part of the adjustment. Large-scale climate responses can involve shifts of probability distributions. To partially account for this, we quantify the response in the variance of the net TOA radiative flux to surface warming perturbations. As above, the estimate is based on an ensemble of $10^{6}$ trajectories. For consistency in physical units, we express the response in terms of standard deviation rather than variance, denoting it by $R^{\sigma}_{\text{TOA},j}(1)$ ($\sigma$ indicates the standard deviation). This quantity measures the change in ensemble spread following a perturbation of the $j$-th SST component.
Conceptually, this procedure provides a computationally efficient analogue of atmospheric-only Green’s function experiments, while shifting the focus from individual trajectories to full probability distributions.
Sensitivity maps.
@@
The vectors of one-time-step responses in mean and standard deviation, $\mathbf{R}^{\mu}$ and $\mathbf{R}^{\sigma}$, are projected back to the full $N$-dimensional SST field using the EOFs, $\mathbf{e}\in\mathbb{R}^{n,N}$:
$\begin{split}\mathbf{S}^{\mu,\sigma}=\mathbf{R^{\mu,\sigma}_{\text{TOA}}}(1)\cdot\mathbf{e}=\sum_{j}^{n}R^{\mu,\sigma}_{\text{TOA},j}(1)~\mathbf{e}_{j},\end{split}$ (13)
@@
where $\mathbf{S}^{\mu},\mathbf{S}^{\sigma}\in\mathbb{R}^{N}$ represent sensitivity maps in mean and standard deviation, respectively. At each ocean grid point $i$, these maps approximate the sensitivities $\frac{\partial x^{\mu,\sigma}_{\text{TOA}}}{\partial x_{\text{SST},i}}$, where $x^{\mu}_{\text{TOA}}$ and $x^{\sigma}_{\text{TOA}}$ respectively represent the ensemble mean and standard deviation of the mean, net radiative flux and $x_{\text{SST},i}$ is the SST at location $i$.
@@
The sensitivity maps in mean ($\mathbf{S}^{\mu}$) and standard deviation ($\mathbf{S}^{\sigma}$) are shown in Figure 7. The obtained sensitivity map $\mathbf{S}^{\mu}$ is in qualitative agreement with previous results in response theory (Figure 2(a) in [102]), with other data-driven approaches (see Figure 7(b) in [111] and Figure 2 in [112]), but, most importantly, with the sensitivity map obtained by directly perturbing the SST boundary condition of atmosphere-only climate models (see third row of Figure 2 in [38], see Figure 5(a) in [39]). In the latter case, the maps exhibit fewer spatial features due to the coarse resolution of the perturbation patches (spanning tens of degrees) in numerical models experiments.
@@
Figure 7: Sensitivity maps obtained using the neural emulator in Eq. (11). At each SST grid point $i$, we show the one–time-step response in the ensemble mean ($\mathbf{S}^{\mu}$) and ensemble standard deviation ($\mathbf{S}^{\sigma}$) of the global-mean net radiative flux at the TOA following an impulse perturbation in SST at that location. The computation is carried out in a 21-dimensional reduced space spanned by empirical orthogonal functions, and the maps represent the projection of these low-dimensional results back to the full grid. Responses are estimated from an ensemble of $10^{6}$ initial conditions.
@@
The dominant feature of $\mathbf{S}^{\mu}$ is a dipole in the tropical Pacific: negative sensitivity in the west and positive sensitivity in the east [39]. This implies that eastern Pacific warming produces a positive radiative feedback, while western Pacific warming produces the opposite. The mechanisms are well established [113, 100]: in the east, high climatological low-level cloud cover decreases under warming, reducing reflected shortwave radiation and yielding positive sensitivity. In the west, deep convection transports surface warming vertically and then horizontally via gravity waves, strengthening the inversion over the eastern Pacific and other stratocumulus regions, increasing low-level cloud fraction and reflected shortwave radiation, thus producing a negative feedback. This analysis confirms that the accurate reconstructions in the forced regime (Fig. 6e,f) arise from physically consistent mechanisms, and that coarse-grained neural models can recover such causal linkages.
Finally, the variance response $\mathbf{S}^{\sigma}$ (Fig. 7b) highlights regions where perturbations alter the spread of radiative responses across ensemble members. Perturbations in stratocumulus-dominated regions (northeast Pacific, off the coast of South America) yield negative variance responses: the ensemble distribution of net radiative flux narrows relative to the control, highlighting small uncertainties in the mean response studied above. By contrast, surface warming perturbations in the central and western Pacific lead to positive variance responses, reflecting a broader spread of the net radiative flux distribution after SST perturbations. To our knowledge, such variance responses have not been quantified before. Unlike the response in the mean, which has been examined extensively and shows broad agreement across atmospheric models [38], the response in ensemble variance has no established benchmark. Because it reflects changes in variability, it may differ significantly across atmospheric models. Moreover, estimating such responses requires very large ensembles, which are prohibitively expensive in conventional climate models. For this reason, the results presented here should be viewed as an exploratory step: they highlight potentially important regional structures, but a full climate interpretation will require systematic intercomparison and validation against other modeling frameworks, such as novel data-driven tools to estimate response through the fluctuation-dissipation theorem [114].
IV.2.4 Main takeaways
@@
In this Section, we developed a stochastic neural emulator that captures the joint evolution of the sea surface temperature (SST) field and the global mean net radiative flux at the top of the atmosphere (TOA) in the GFDL-CM4 climate model. Addressing a specific, well-posed problem, proved crucial in guiding the construction of the model: it enabled us to build on previous literature, restrict attention to just two variables and a limited range of spatial and temporal scales. These simplifications led to a stable and computationally efficient emulator that responds to external forcing and can be used to investigate causal relationships via perturbation experiments. Such experiments focused on very large ensembles ($10^{6}$ trajectories), currently impossible to perform with climate models. These results demonstrate that reduced-order neural emulators can extend beyond reproducing stationary climate dynamics to enable efficient, ensemble-based perturbation experiments.
V Discussion and Conclusion
@@
In summary, the first part of this work discussed the limitations of purely data-driven models of complex climate systems. We highlighted the potential of combining traditional reduced-order modeling approaches, including state-dependent (multiplicative) stochastic parameterizations, with recent advances in neural networks to study large-scale, low-frequency climate dynamics. We emphasized the need to evaluate such models against both stationary and perturbed statistics. In this context we discussed how response theory provides rigorous guidance to move beyond stationary evaluation and probe causal mechanisms. These ideas were then illustrated through a climate-dynamics problem focusing on inferring causal relations between surface warming and changes in net radiative fluxes. Given careful variable selection, reduced-order neural stochastic models can capture both stationary and perturbed statistics in a computationally efficient way. Below, we expand on these points and outline directions for future work.
@@ V.1 Limits of data-driven-only climate emulators and the role of stochastic approaches
    Unknown equations, partially observed state vector. This more realistic case presents fundamental difficulties, even with access to high-quality data and powerful neural models architectures. The model skill now depends critically on two factors: (a) the choice of variables to model and (b) how unobserved degrees of freedom are parameterized.
These results highlight intrinsic limitations of purely data-driven-only modeling in high-dimensional climate systems, where the full set of relevant variables is typically unknown and no rigorous framework exists for selecting the proper variables to model. By variable selection we mean the choice of climate fields as well as their spatial and temporal scales (also including dimensionality reduction methodologies). Such choices are guided by the physical process of interest, and in the case of low-frequency climate dynamics, they motivate the development of reduced-order neural emulators tailored to specific phenomena as efficient alternative to general purpose emulators. A central challenge is to parameterize the cumulative effect of unresolved degrees of freedom. Model performance (especially in capturing responses to external forcings) depends critically on this choice and we showed that multiplicative (state-dependent) noise is preferable to additive noise.
We emphasize that the analysis presented here applies to data-driven models of turbulent geophysical systems that must capture both stationary statistics and responses to external perturbations, as in (causal) Green’s function experiments [38]. For Earth system models, these issues mainly concern autoregressive climate emulators, particularly coupled atmosphere–ocean systems, and should not be confused with short-term weather forecasting using neural networks (e.g., [115, 23]), where long-term stability and forced responses are less relevant. Our study also does not address statistics-based emulators for climate change projections, which have proven effective for investigating responses of key climate variables to CO2 forcing [116, 117]; see [116] for an example where simple approaches, such as linear pattern scaling, can outperform more complex models.
@@
The discussion above suggests two possible directions for future research in autoregressive emulators of climate. First, a stochastic formulation of neural autoregressive climate emulators is preferable to a purely deterministic one. Probabilistic formulations align more closely with the inherently stochastic and uncertain nature of climate dynamics [118]. Recent advances in weather emulation already point in this direction such as FourecastNet3 [23], GenCast [23] and the novel work of Alet et al. (2025) [119]. Second, the ability of climate emulators in capturing forced responses in the full probability distribution, akin to Green’s function experiments, remains largely unexplored. A promising way forward is to evaluate novel architectures on their ability to respond to external perturbations on a hierarchy of simplified models relevant to geophysical fluid dynamics, such as the Lorenz–96 system [120]. These controlled testbeds enable rigorous quantification of forced responses and could inform the design of future architectures for data-driven climate modeling.
@@ V.2 Response theory as a theoretical framework for model evaluation
We argued that the impulse response operator is a building block that a model should capture in order to reproduce causal interactions across degrees of freedom and to accurately respond to external perturbations (Section II.2 and Appendix A). This perspective enables us to move beyond evaluation metrics based solely on stationary statistics and instead consider the effect of external interventions. Assessing impulse responses is straightforward in the case of data-driven emulators built from simple numerical models, i.e. models that allow us to simulate thousands of ensemble members, as in the first part of this study. In such cases, one can actively perturb the reference model to generate a ground truth for the response operator. However, for models constructed from observational data, such direct perturbations are not possible. In this context, one could, in principle, leverage the Fluctuation-Dissipation Theorem (FDT) to predict the response operator from data alone [81, 17]. Traditionally, the quasi-Gaussian approximation have yield reliable estimates of impulse responses in the ensemble mean [121, 47]. Recent advances in score-based generative models have shown the potential to extend these predictions to responses in higher-order moments [114].
@@
In the absence of direct perturbation experiments, FDT provides a rigorous bridge between stationary and forced dynamics [49, 48]. With sufficiently long integrations of current climate emulators, one can readily focus on representative large-scale modes and estimate their impulse responses from the stationary record, thereby extending evaluation beyond stationary statistics with minimal computational cost. A recent example illustrating this approach in the climate context can be found in [18]. Moreover, the response operator at the shortest time step encodes direct causal links, i.e. the input–output relationships across variables in a reduced-order emulator (see Appendix A). Future work will leverage this information to suppress spurious, non-causal input-output dependencies during training, similarly to the recent work of Chen and Liu (2024) [59] on causally constrained, sparse reduced-order models.

### v4

@@ IV.2.3 Interpretability and causal understanding through ensemble mean responses
In the previous section, we showed that the reduced-order neural emulator can reconstruct the net radiative flux at the top of the atmosphere (TOA) from prescribed changes in the sea surface temperature (SST) field. This amounts to the mapping $\mathbf{x}_{\text{SST}}^{\text{f}}(t)\rightarrow x_{\text{TOA}}(t+1)$, where $\mathbf{x}^{\text{f}}_{\text{SST}}(t)\in\mathbb{R}^{n}$, $x_{\text{TOA}}(t)\in\mathbb{R}$, and $x_{\text{TOA}}(0)=0$. We now ask the following question: What are the SST patterns that contribute the most to this skillfull prediction? Addressing this question requires moving beyond prediction tasks and identify the causal influence of individual SST components on radiative fluxes. We frame the problem in terms of linear response theory, focusing on the direct link $x_{j,\text{SST}}(t)\rightarrow x_{\text{TOA}}(t+1)$, where $x_{j,\text{SST}}(t)$ denotes the $j$-th SST component. We refer the reader to Appendix A for details on the causal interpretation of the response operator.
We apply an impulse to the $j$-th SST component at $t=0$ and compute the ensemble-mean response of the net TOA flux at $t=1$, denoted $\delta\langle x_{\text{TOA}}(1)\rangle$ (see Appendix A for details). This quantity, referred to as $R_{\text{TOA},j}(1)$, represents the one-step ensemble-mean response and is estimated using $10^{6}$ initial conditions to ensure robust statistics. By focusing on the one-step response, we isolate fast atmospheric adjustments while neglecting the slower oceanic feedbacks (see Section 7 in [FDTPatternEffect] for details). This setup is thus conceptually closer to atmospheric-only Green’s function experiments, where the ocean acts as a prescribed boundary condition rather than a dynamically coupled component of the system.
@@
The vector $\mathbf{R}_{\text{TOA}}(1)\in\mathbb{R}^{n}$ is projected back to the full $N$-dimensional SST field using the EOFs, $\mathbf{e}\in\mathbb{R}^{n,N}$:
$\begin{split}\mathbf{S}=\mathbf{R_{\text{TOA}}}(1)\cdot\mathbf{e}=\sum_{j}^{n}R_{\text{TOA},j}(1)~\mathbf{e}_{j}.\end{split}$ (13)
@@
At each ocean grid point $i$, the sensitivity map $\mathbf{S}\in\mathbb{R}^{N}$ approximates the sensitivity $\frac{\partial x_{\text{TOA}}}{\partial x_{\text{SST},i}}$, where $x_{\text{TOA}}$ is the global mean radiative flux at the TOA and $x_{\text{SST},i}$ is the SST at location $i$.
@@
The sensitivity map $\mathbf{S}$ is shown in Figure 7. The obtained sensitivity map $\mathbf{S}$ is in qualitative agreement with previous results in response theory (Figure 2(a) in [FDTPatternEffect]), with other data-driven approaches (see Figure 7(b) in [kang2023recent] and Figure 2 in [senneNN]), and with the one obtained by directly perturbing the SST boundary condition of atmosphere-only climate models (see third row of Figure 2 in [blochjohnsonetal2024], see Figure 5(a) in [Dong2019AttributingPacific]). In the latter case, the maps exhibit fewer spatial features due to the coarse resolution of the perturbation patches (spanning tens of degrees) in numerical models experiments. For additional statistical reconstruction approaches to the pattern effect that yield structures similar to those in Figure 7, see the recent analysis by Fredericks et al. (2025) [Leif].
@@
Figure 7: Sensitivity maps obtained using the neural emulator in Eq. (11). At each SST grid point $i$, we show the one–time-step response in the ensemble mean ($\mathbf{S}^{\mu}$) of the global-mean net radiative flux at the TOA following an impulse perturbation in SST at that location. The computation is carried out in a 21-dimensional reduced space spanned by empirical orthogonal functions, and the maps represent the projection of these low-dimensional results back to the full grid.
@@
The dominant feature of $\mathbf{S}$ is a dipole in the tropical Pacific: negative sensitivity in the west and positive sensitivity in the east [Dong2019AttributingPacific]. This implies that eastern Pacific warming produces a positive radiative feedback, while western Pacific warming produces the opposite. The mechanisms are well established [Pascale, williams2023circus]: in the east, high climatological low-level cloud cover decreases under warming, reducing reflected shortwave radiation and yielding positive sensitivity. In the west, deep convection transports surface warming vertically and then horizontally via gravity waves, strengthening the inversion over the eastern Pacific and other stratocumulus regions, increasing low-level cloud fraction and reflected shortwave radiation, thus producing a negative feedback. This analysis confirms that the accurate reconstructions in the forced regime (Figure 6(e,f)) arise from physically consistent mechanisms.
IV.2.4 Limitations and a few ways forward
Our coarse-grained representation enabled the development of a data-driven model to study the direct causal interaction $x_{\text{SST}}\rightarrow x_{\text{TOA}}$ through perturbation experiments, while also capturing the stationary distribution well. However, these coarse-graining choices also introduce some drawbacks. We discuss these limitations below, further delineating the model’s domain of validity and outlining possible ways forward:
  • The model does not fully capture the decay of autocorrelations. For example, the first SST mode in Figure 6(a) suggests residual non-Markovian behavior. This is not an issue for our focus on direct causal links (responses at the shortest time step), but it would matter for responses over longer time horizons. Remedies include: (i) expanding the state vector by including slower oceanic variables and more EOF components, (ii) incorporating memory terms in the stochastic parameterization [KONDRASHOV201533], or (iii) further coarse-graining the temporal resolution (e.g., 3-month averaging). Such choices will depend on the intended scientific application.
  • We found that responses in ensemble variance are sensitive to retraining the multiplicative noise term, so we do not present these results. The fitting algorithm in Appendix C requires far more data than available here (only 7200 samples for training and validation), to achieve high accuracy. In this data-scarce setting, exploring simple analytical forms for the multiplicative noise may be beneficial and it is left for future work. Importantly, data-driven approaches, such as the Fluctuation–Dissipation Theorem (FDT) [MajdaBook, Marconi], can provide an approximate benchmark for responses. This allows calibration of models with different stochastic terms, particularly to test variance responses where direct benchmarks are often unavailable. This strategy aligns with recent proposals such as [majdaSIAM] and will be explored in future work.
  • While our emulators are empirically stable, stability is not guaranteed, which may matter if tested against large perturbations. Future work could impose additional constraints on the neural architectures, similar to previous approaches for physically constrained regression models as in [majdaConstraints, KONDRASHOV201533].
IV.2.5 Main takeaways
@@
In this Section, we developed a stochastic neural emulator that captures the joint evolution of the sea surface temperature field and the global mean net radiative flux at the top of the atmosphere in the GFDL-CM4 climate model. Addressing a specific, well-posed problem, proved crucial in guiding the construction of the model: it enabled us to build on previous literature, restrict attention to just two variables and a limited range of spatial and temporal scales. These simplifications led to a stable and computationally efficient emulator that responds to external forcing and can be used to investigate causal relationships via perturbation experiments. Such experiments focused on very large ensembles ($10^{6}$ trajectories), currently impossible to perform with climate models.
V Discussion
@@
In summary, the first part of this work discussed important limitations of purely data-driven models of complex climate systems. We highlighted the potential of combining traditional reduced-order modeling approaches, including state-dependent (multiplicative) stochastic parameterizations, with recent advances in neural networks to study large-scale, low-frequency climate dynamics. We emphasized the need to evaluate such models against both stationary and perturbed statistics. In this context we discussed how response theory provides rigorous guidance to move beyond stationary evaluation and probe causal mechanisms. These ideas were then illustrated through a climate-dynamics problem focusing on inferring causal relations between surface warming and changes in net radiative fluxes. Given careful variable selection, reduced-order neural stochastic models can capture both stationary and perturbed statistics in a computationally efficient way. Below, we expand on these points and outline directions for future work.
@@ V.1 Limits of data-driven-only climate emulators and the role of stochastic approaches
    Unknown equations, partially observed state vector. This more realistic case presents fundamental difficulties, even with access to high-quality data and powerful neural model architectures. The most critical limitation is (a) identifying an appropriate coarse-grained representation; once this is determined, (b) the cumulative effect of unobserved degrees of freedom needs to be accurately parametrized. Learning such parametrizations from data remains a central challenge, and we showed that multiplicative noise can be advantageous over additive noise, though it requires a larger dataset for reliable training.
These results highlight intrinsic limitations of purely data-driven modeling of high-dimensional, multiscale climate systems, where the full set of relevant variables is unobserved or even unknown  [^2] A brief remark: in partially observed settings, one might be tempted to appeal to Takens’ embedding theorem [Takens]. This is not a practical option: the theorem does not extend to stochastic systems, and it becomes infeasible in high-dimensional settings [Cecconi, Baldovin, Lucente].. In this context, the “correct” choice of variables, scales, and stochastic parameterizations depends ultimately on the specific scientific goals [Smith2001]. This perspective helps explain the continued success of Linear Inverse Models (LIMs) proposed by Penland, C. (1989) [Penland89], which rely on careful selection of climate fields and preprocessing choices to capture targeted aspects of climate variability; see, for example, [Penland95, PENLAND1996534, Penland2025] among others.
We emphasize that the analysis presented here applies to data-driven models of turbulent geophysical systems that must capture both stationary statistics and responses to external perturbations, as in (causal) Green’s function experiments [blochjohnsonetal2024]. For Earth system models, these issues mainly concern autoregressive climate emulators, particularly coupled atmosphere–ocean systems, and should not be confused with short-term weather forecasting using neural networks (e.g., [AIFS, fourcastnet3]), where long-term stability and forced responses are less relevant. Our study also does not address statistics-based emulators for climate change projections, which have proven effective for investigating responses of key climate variables to CO2 forcing [Bjorn, Gosha, AndreEmulation].
@@
The discussion above suggests two possible directions for future research in autoregressive emulators of climate. First, a stochastic formulation of neural autoregressive climate emulators is preferable to a purely deterministic one. Probabilistic formulations align more closely with the inherently stochastic and uncertain nature of climate dynamics [PalmerQJRMS]. Recent advances in weather emulation already point in this direction such as FourecastNet3 [fourcastnet3], GenCast [fourcastnet3] and the novel work of Alet et al. (2025) [Alet]. New interesting directions also include probabilistic emulation in the framework of stochastic interpolants [ProbabilisticSI, ProbabilisticSIPedram, Nikolaj]. Second, the ability of climate emulators in capturing forced responses in the full probability distribution, akin to Green’s function experiments, remains largely unexplored, especially for responses in higher-order moments. A promising way forward is to evaluate novel architectures on their ability to capture the mean and variance response to external perturbations in a hierarchy of simplified models relevant for large scale climate variability, such as the 2-scale Lorenz–96 system [Lorenz96]. These controlled testbeds enable rigorous quantification of forced responses and could inform the design of future architectures for data-driven climate modeling.
@@ V.2 Response theory as a theoretical framework for model evaluation
We argued that the impulse response operator is a building block that a model should capture in order to reproduce causal interactions across degrees of freedom and to accurately respond to external perturbations (Section II.2 and Appendix A). This perspective enables us to move beyond evaluation metrics based solely on stationary statistics and instead consider the effect of external interventions. Assessing impulse responses is straightforward in the case of data-driven emulators built from simple numerical models, i.e. models that allow us to simulate thousands of ensemble members, as in the first part of this study. In such cases, one can actively perturb the reference model to generate a ground truth for the response operator. However, for models constructed from observational data, such direct perturbations are not possible. In this context, one can leverage the Fluctuation-Dissipation Theorem (FDT) to predict the response operator from data alone [MajdaStructuralStability, Baldovin]. Traditionally, the quasi-Gaussian approximation have yield reliable estimates of impulse responses in the ensemble mean [GRITSUN, Majda2010]. Recent advances in score-based generative models have shown the potential to extend these predictions to responses in higher-order moments [LudovicoFabriAndre].
@@
In the absence of direct perturbation experiments, FDT provides a rigorous bridge between stationary and forced dynamics [Marconi, MajdaBook]. With sufficiently long integrations of current climate emulators, one can readily focus on representative large-scale modes and estimate their impulse responses from the stationary record, thereby extending evaluation beyond stationary statistics at minimal computational cost. A recent example illustrating this approach in the climate context can be found in [FabCausal]. Importantly, the analysis reveals the need to test emulators in their ability to respond to higher-order moments. [LudovicoFabriAndre] have shown that coupling the FDT with generative models provide a flexible framework to estimate responses in the full probability distribution. Moreover, the response operator at the shortest time step encodes direct causal links, i.e. the input–output relationships across variables in a reduced-order emulator (see Appendix A). Future work will leverage this information to suppress spurious, non-causal input-output dependencies during training, similarly to the recent work of Chen and Liu (2024) [nanChen], leading to causally constrained, sparse reduced-order models.